# Guided Transfer Learning for Discrete Diffusion Models

### v2

@@ 1 Introduction
Outside of transfer learning, Schiff et al. (2025); Nisonoff et al. (2025) have explored CBG mechanisms for discrete DMs. These methods apply mainly to domains with small $|\mathcal{V}|$ and modest $L$, such as biological sequences, leaving scalable CBG for language modeling as an open problem. This constaints are due to the necessity of token substitutions across all positions and vocabulary items during score computation. More precisely, letting $L$ be the sequence length and $\mathcal{V}$ the vocabulary, naive CBG requires $\mathcal{O}(L|\mathcal{V}|)$ guidance evaluations per step, which is infeasible for long sequences and large vocabularies. In this paper, we aim to leverage CBGs for transfer learning, while overcoming the limiting small vocabulary sizes.
@@ 2 Background
Diffusion models are generative models defined by a forward noising process and a learned reverse denoising process. In the reverse process, a denoiser $p_{\theta}$ is trained to remove noise from latent variables $\mathbf{z}_{t}$ to recover data $x$. The forward process generates $\mathbf{z}_{t}$ by injecting noise into clean data. We consider discrete latents: $\mathbf{z}_{t}\in\mathcal{V}$, the set of one-hot tokens over a vocabulary of size $N$, $\mathcal{V}\coloneqq\{\mathbf{z}\in\{0,1\}^{N}:\sum_{i=1}^{N}z_{i}=1\}\subset\Delta^{N-1},$ where $\Delta^{N-1}$ denotes the probability simplex over $N$ categories. Let $\text{Cat}(\cdot;\mathbf{\pi})$ be the categorical distribution over $N$ classes with the probability vector $\mathbf{\pi}\in\Delta^{N-1}$.
@@ 3 Guided Transfer Learning for DDMs
Practical implementation. Equation 6 (proved in Appendix C) expresses the target reverse transition as a reweighted version of the source reverse transition. This expression suggests a direct implementation with two components: a frozen source denoiser that provides $p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})$ and a learned ratio module that provides the ratio-dependent weight. Concretely, we approximate the source kernel with the pretrained denoiser $p_{\theta}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})$, and we learn a ratio function $r_{\phi}(\mathbf{z}_{s(i)})\approx\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\!\left[\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\right].$ Together, these two networks define the GTL sampler, $\psi^{\star}=[\theta,\phi]$, which generates target-domain samples while keeping the denoiser fixed.
@@ 3.1 Scalability via Planner Sampling
Caching. In masked diffusion, the timestep is implicitly determined by the number of masked tokens, allowing us to remove explicit time conditioning from the networks (Sahoo et al., 2024). As a result, if a denoising step does not modify the sequence, the corresponding logits can be reused. By caching logits across steps, the number of network evaluations becomes upper-bounded by the number of unmasking events. In the worst case, each position unmasks once, yielding at most $L$ effective network calls. This optimization does not apply to uniform-noise diffusion, where token identities are mixed. Top-$n_{\text{ratio}}$ pruning. Although caching reduces the number of steps, each step still requires evaluating the ratio network over the full vocabulary. To address this, we restrict ratio evaluations to a small candidate set. We assume that the denoiser distribution $p_{\theta}(\cdot\mid\mathbf{z}_{t})$ is sufficiently concentrated, so that meaningful deviations between the source and target distributions occur primarily among the top-$n_{\text{ratio}}$ tokens. Formally, we assume that for these candidates the density ratio $r(z)=q(z)/p(z)$ is bounded by a constant $C<\infty$; no such bound is required outside this set, where $p(z)$ is already negligible. Under this assumption, we evaluate the ratio term only for the top-$n_{\text{ratio}}$ tokens proposed by the denoiser at each position. This produces position-specific candidate sets in the denominator of equation 7 and reduces the cost of ratio evaluations to $\mathcal{O}(n_{\text{ratio}}\,L)$, with $n_{\text{ratio}}\ll|\mathcal{V}|$ in practice.
Mask-probability stabilizer. The interaction between top-$n_{\text{ratio}}$ pruning and the guidance weight $\gamma$ can distort probability mass, occasionally causing the diffusion process to collapse to fully masked sequences. To prevent this failure mode, we enforce a position-independent constraint that preserves the probability of remaining masked across source and target distributions. This induces a renormalized update over non-mask tokens and stabilizes the denoising dynamics. Details of this stabilization are provided in the appendix.
@@
Even with top-$n_{\text{ratio}}$ pruning, the overall complexity still scales linearly with $L$. To further reduce cost, we introduce a planner-based sampler that selects a single position to update at each step. The planner $\rho_{\vartheta}$ is trained to predict the next denoising position in the masked sequence, replacing the stochastic Gillespie process with a deterministic $\tau$-leaping schedule. This strategy fixes the number of denoising steps to exactly $L$, which is comparable to standard masked diffusion where typically $T\approx L$, while substantially reducing guidance evaluations and wall-clock time. Algorithm 1 summarizes the resulting ratio-guided denoising step. Crucially, once the planner selects a position, ratio evaluations over the top-$n_{\text{ratio}}$ candidates can be fully vectorized. As a result, each denoising step requires only a single forward pass through the denoiser, ratio network, and planner. For planner training, we follow simplified strategies from Peng et al. (2025); Liu et al. (2025). The ratio network is trained using a discrete-data loss adapted from Ouyang et al. (2024). Additional training details are provided in the supplementary materials.
@@ 4.1 Synthetic Dataset
n Vanilla Discrete Diffusion Finetuned Discrete Diffusion GTL
@@ 20 Refer to caption Refer to caption Refer to caption
Figure 2: True source and target (left). Estimated target transition matrices for varying $n$ across three methods (right).
Experimental Setup
We construct two Markov chains with distinct transition matrices. Let the probability state vector be $\mathbf{\pi}_{i}\in\Delta^{N-1}$ over a vocabulary of size $N=5$. The transition matrix of the target domain is $\mathbf{T}\in\mathbb{R}^{N\times N}$ and that of the source domain is $\mathbf{S}\in\mathbb{R}^{N\times N}$. We sample sequences of length $20$ from the target domain using the update rule $\mathbf{\pi}_{i}=\mathbf{T}\,\mathbf{\pi}_{i-1}$ with the initial state $\mathbf{\pi}_{0}$ set to the uniform vector. Sampling from the source domain proceeds in the same way. Implementation details can be found in the appendix E.1.
Results
We consider $m=10000$ source samples and $n\in\{1000,100,20\}$ target samples. We estimate the transition matrix by sampling $4096$ sequences, counting transitions between states, and normalizing by the total count. Figure 2 presents an example where $\mathbf{S}$ has diagonal entries $0.1$ and $\mathbf{T}$ has diagonal entries $0.8$. We observe that the vanilla target diffusion model and the finetuned model both degrade in performance when fewer target samples are available.
$n_{\text{tgt}}$ Vanilla (Target-only) Finetuned GTL
@@ Results
Table 2: Comparison of baseline methods and best guided results (mean $\pm$ standard error; n=3n{=}3 seeds) across target sample sizes for $(\text{diag}_{\text{src}}=0.1,\text{diag}_{\text{tgt}}=0.8)$. Reported values are the average Kullback–Leibler divergence between true and estimated transition matrices; lower is better.
@@ 4.2 Language Modeling
Refer to caption Refer to caption Refer to caption
(a) Effect of $n_{\text{ratio}}$ at $\gamma=5$. (b) Effect of low $\gamma$ at fixed $n_{\text{ratio}}$ (c) MAUVE vs. high $\gamma$
@@
In this section, we apply our method to text data. We use the arXiv abstracts dataset (arXiv.org submitters, 2024), which contains 1.7 million arXiv articles with titles, authors, categories, and abstracts. We focus on Computer Science, Mathematics, and Physics. For data preparation, we concatenate all text and split it into segments of 512 tokens using the bert-base-uncased tokenizer with vocabulary size N=30,522N=30{,}522. This results in 285,946285{,}946 samples in Computer Science, 176,831176{,}831 in Mathematics, and 79,63179{,}631 in Physics. For each category we randomly select 10% of the samples for validation. We ensure that the three category subsets are disjoint. We define the union of Computer Science and Mathematics as the source domain and Physics as the target domain.
@@
Implementation and Evaluations
We train the source diffusion model for 100,000 gradient steps using the DiT architecture (Peebles and Xie, 2023) with 59.8M trainable parameters. A log–linear noise scheduler is used for both training and sampling. The finetuned model on the target domain is trained for 10,000 gradient steps. For our guided method, we initialize a ratio model with 4.1M parameters and train it for 5,000 gradient steps. Additional training details appear in the supplementary materials.
For each method we generate 128 sequences of length 512. We evaluate quality with MAUVE (Pillutla et al., 2021), which captures the tradeoff between perplexity and diversity. Although generative perplexity (Gen. PPL) is often reported, we observed, consistent with prior work (Wang et al., 2025), that Gen. PPL can remain low even when the text has reduced utility, such as repeated words. In those cases, MAUVE is more sensitive and provides a clearer measure of quality. We also evaluate the domain of each sampled sequence using an independently trained classifier. Scores near zero indicate the target domain and scores near one indicate the source domain. The classifier follows the setup used for the ratio model in the supplementary materials with label smoothing set to $0.05$.
@@ 4.2.2 Method Comparison and Sensitivity to Data Fraction
We compare three methods: a fine-tuned discrete diffusion model trained on the target domain after pretraining on the source domain, a vanilla discrete diffusion model trained directly on the target domain, and our Guided Transfer Learning (GTL) method with the ratio estimator trained across target-data fractions of 100%, 30%, 5%, and 1%. Each method is trained with three random seeds, and sampling is performed with three seeds as well (nine runs per cell). Furthermore, each seed corresponds to a distinct subset of the target data. Figure 1 reports the behavior of the models across different fractions of the target training data. The observed trends align with those in Table 2. In the smallest case, we only use 796 samples of Physics abstracts. For finetuned and vanilla discrete diffusion, the number of epochs over the training set increases up to 400 when the fraction is smallest. The ratio model decreases its MAUVE score with decreasing fractions, but less dramatically than the other methods. Thus, our method demonstrates greater stability under limited target-domain data. In contrast, the finetuned and vanilla discrete diffusion models show a stepwise drop in performance as the fraction decreases. The GTL remains relatively stable, except in the 1% case, where we also observe a sharp decline.

### v4

@@ 1 Introduction
Outside of transfer learning, Schiff et al. (2025); Nisonoff et al. (2025) have explored CBG mechanisms for discrete DMs. These methods apply mainly to domains with small vocabulary size $|\mathcal{V}|$ and modest sequence length $L$, such as biological sequences, leaving scalable CBG for language modeling as an open problem. This constaints are due to the necessity of token substitutions across all positions and vocabulary items during score computation. More precisely, naive CBG requires $\mathcal{O}(L|\mathcal{V}|)$ guidance evaluations per step, which is infeasible for long sequences and large vocabularies. In this paper, we aim to leverage CBGs for transfer learning, while overcoming the limiting small vocabulary sizes.
@@ 2 Background
Diffusion models are generative models defined by a forward noising process and a learned reverse denoising process. In the reverse process, a denoiser $p_{\theta}$ is trained to remove noise from latent variables $\mathbf{z}_{t}$ to recover data $x$. The forward process generates $\mathbf{z}_{t}$ by injecting noise into clean data. We consider discrete latents: $\mathbf{z}_{t}\in\mathcal{V}$, the set of one-hot tokens over a vocabulary $\mathcal{V}$ of size $N$, $\mathcal{V}\coloneqq\{\mathbf{z}\in\{0,1\}^{N}:\sum_{i=1}^{N}z_{i}=1\}\subset\Delta^{N-1},$ where $\Delta^{N-1}$ denotes the probability simplex over $N$ categories. Let $\text{Cat}(\cdot;\mathbf{\pi})$ be the categorical distribution over $N$ classes with the probability vector $\mathbf{\pi}\in\Delta^{N-1}$.
@@ 3 Guided Transfer Learning for DDMs
Practical implementation. Equation 6 (proved in Appendix C) expresses the target reverse transition as a reweighted version of the source reverse transition. This expression suggests a direct implementation with two components: a frozen source denoiser that provides $p(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})$ and a learned ratio module that provides the ratio-dependent weight. Concretely, we approximate the source kernel with the pretrained denoiser $p_{\theta}(\mathbf{z}_{s(i)}\mid\mathbf{z}_{t(i)})$, and we learn a ratio function $r_{\phi}(\mathbf{z}_{s(i)})\approx\mathbb{E}_{\mathbf{x}_{0}\sim p(\cdot\mid\mathbf{z}_{s(i)})}\!\left[\frac{q(\mathbf{x}_{0})}{p(\mathbf{x}_{0})}\right].$ Together, these two networks define the GTL sampler, $\psi^{\star}=[\theta,\phi]$, which generates target-domain samples while keeping the denoiser fixed. The ratio network is trained using a discrete-data loss adapted from Ouyang et al. (2024). Additional training details are provided in the supplementary materials.
@@ 3.1 Scalability via Planner Sampling
Caching. In masked diffusion, the timestep is implicitly determined by the number of masked tokens, allowing us to remove explicit time conditioning from the networks (Sahoo et al., 2024). As a result, if a denoising step does not modify the sequence, the corresponding logits can be reused. By caching logits across steps, the number of network evaluations becomes upper-bounded by the number of unmasking events. In the worst case, each position unmasks once, yielding at most $L$ effective network calls. This optimization does not apply to uniform-noise diffusion, where token identities are mixed. Top-$n_{\text{ratio}}$ pruning. Although caching reduces the number of steps, each step still requires evaluating the ratio network over the full vocabulary. To address this, we restrict ratio evaluations to a small candidate set. We assume that the denoiser distribution $p_{\theta}(\cdot\mid\mathbf{z}_{t})$ is sufficiently concentrated, so that meaningful deviations between the source and target distributions occur primarily among the top-$n_{\text{ratio}}$ tokens. Formally, we assume that for these candidates the density ratio $r(z)=q(z)/p(z)$ is bounded by a constant $C<\infty$; no such bound is required outside this set, where $p(z)$ is already negligible. Under this assumption, we evaluate the ratio term only for the top-$n_{\text{ratio}}$ tokens proposed by the denoiser at each position. This produces position-specific candidate sets in the denominator of equation 7 and reduces the cost of ratio evaluations to $\mathcal{O}(n_{\text{ratio}}\,L)$, with $n_{\text{ratio}}\ll|\mathcal{V}|$ in practice. Mask-probability stabilizer. The interaction between top-$n_{\text{ratio}}$ pruning and the guidance weight $\gamma$ can distort probability mass, occasionally causing the diffusion process to collapse to fully masked sequences. To prevent this failure mode, we enforce a position-independent constraint that preserves the probability of remaining masked across source and target distributions. This induces a renormalized update over non-mask tokens and stabilizes the denoising dynamics. Details of this stabilization are provided in the appendix.
@@
Even with top-$n_{\text{ratio}}$ pruning, the overall complexity still scales linearly with $L$. To further reduce cost, we introduce a planner-based sampler that selects a single position to update at each step. The planner $\rho_{\vartheta}$ is trained to predict the next denoising position in the masked sequence, replacing the stochastic Gillespie process with a deterministic $\tau$-leaping schedule. This strategy fixes the number of denoising steps to exactly $L$, which is comparable to standard masked diffusion where typically $T\approx L$, while substantially reducing guidance evaluations and wall-clock time. Algorithm 1 summarizes the resulting ratio-guided denoising step. Crucially, once the planner selects a position, ratio evaluations over the top-$n_{\text{ratio}}$ candidates can be fully vectorized. As a result, each denoising step requires only a single forward pass through the denoiser, ratio network, and planner. For planner training, we follow simplified strategies from Peng et al. (2025); Liu et al. (2025).
@@ 4.1 Synthetic Dataset
$\bm{n}$ Vanilla Discrete Diffusion Finetuned Discrete Diffusion GTL
@@ 20 Refer to caption Refer to caption Refer to caption
Figure 2: True source and target (left). Estimated target transition matrices for varying number of target samples $n$ across three methods (right).
Experimental Setup We construct two Markov chains with distinct transition matrices. Let the probability state vector be $\mathbf{\pi}_{i}\in\Delta^{N-1}$ over a vocabulary of size $N=5$. The transition matrix of the target domain is $\mathbf{T}\in\mathbb{R}^{N\times N}$ and that of the source domain is $\mathbf{S}\in\mathbb{R}^{N\times N}$. We sample sequences of length $20$ from the target domain using the update rule $\mathbf{\pi}_{i}=\mathbf{T}\,\mathbf{\pi}_{i-1}$ with the initial state $\mathbf{\pi}_{0}$ set to the uniform vector. Sampling from the source domain proceeds in the same way. Implementation details can be found in the appendix E.1.
Results We consider $m=10000$ source samples and $n\in\{1000,100,20\}$ target samples. We estimate the transition matrix by sampling $4096$ sequences, counting transitions between states, and normalizing by the total count. Figure 2 presents an example where $\mathbf{S}$ has diagonal entries $0.1$ and $\mathbf{T}$ has diagonal entries $0.8$. We observe that the vanilla target diffusion model and the finetuned model both degrade in performance when fewer target samples are available.
$n$ Vanilla (Target-only) Finetuned GTL
@@
Table 2: Comparison of baseline methods and best guided results (mean $\pm$ standard error; 33 seeds) across target sample sizes for $(\text{diag}_{\text{src}}=0.1,\text{diag}_{\text{tgt}}=0.8)$. Reported values are the average Kullback–Leibler divergence between true and estimated transition matrices; lower is better.
@@ 4.2 Language Modeling
In this section, we apply our method to text data. We use the arXiv abstracts dataset (arXiv.org submitters, 2024), which contains 1.7 million arXiv articles with titles, authors, categories, and abstracts. We focus on Computer Science, Mathematics, and Physics. For data preparation, we concatenate all text and split it into segments of 512 tokens using the bert-base-uncased tokenizer with vocabulary size N=30,522N=30{,}522. This results in 285,946285{,}946 samples in Computer Science, 176,831176{,}831 in Mathematics, and 79,63179{,}631 in Physics. For each category we randomly select 10% of the samples for validation. We ensure that the three category subsets are disjoint.
@@
Implementation and Evaluations We train the source diffusion model for 100,000 gradient steps using the DiT architecture (Peebles and Xie, 2023) with 59.8M trainable parameters. A log–linear noise scheduler is used for both training and sampling. The finetuned model on the target domain is trained for 10,000 gradient steps. For our guided method, we initialize a ratio model with 4.1M parameters and train it for 5,000 gradient steps. Additional training details appear in the supplementary materials. For each method we generate 128 sequences of length 512. We evaluate quality with MAUVE (Pillutla et al., 2021), which captures the tradeoff between perplexity and diversity. Although generative perplexity (Gen. PPL) is often reported, we observed, consistent with prior work (Wang et al., 2025), that Gen. PPL can remain low even when the text has reduced utility, such as repeated words. In those cases, MAUVE is more sensitive and provides a clearer measure of quality. We also evaluate the domain of each sampled sequence using an independently trained classifier. Scores near zero indicate the target domain and scores near one indicate the source domain. The classifier follows the setup used for the ratio model in the supplementary materials with label smoothing set to $0.05$.
@@ 4.2.2 Method Comparison and Sensitivity to Data Fraction
In this section, we define the union of Computer Science and Mathematics abstracts as the source domain and Physics abstracts as the target domain. We compare three methods: a fine-tuned discrete diffusion model trained on the target domain after pretraining on the source domain, a vanilla discrete diffusion model trained directly on the target domain, and our Guided Transfer Learning (GTL) method with the ratio estimator trained across target-data fractions of 100%, 30%, 5%, and 1%. Each method is trained with three random seeds, and sampling is performed with three seeds as well (nine runs per cell). Furthermore, each seed corresponds to a distinct subset of the target data. Figure 1 reports the behavior of the models across different fractions of the target training data. The observed trends align with those in Table 2. In the smallest case, we only use 796 samples of Physics abstracts. For finetuned and vanilla discrete diffusion, the number of epochs over the training set increases up to 400 when the fraction is smallest. The ratio model decreases its MAUVE score with decreasing fractions, but less dramatically than the other methods. Thus, our method demonstrates greater stability under limited target-domain data. In contrast, the finetuned and vanilla discrete diffusion models show a stepwise drop in performance as the fraction decreases. Even in the full-target-data setting, GTL outperforms the baselines because there is still a large imbalance between the number of source-domain and target-domain samples. The GTL remains relatively stable, except in the 1% case, where we also observe a sharp decline.